If I were to give $ 7 to each of the beggars at my door, I would have $24 left. I lack $32 of being able to give them $9 apiece. What is the sum of my money and the number of beggars at my door ?
Found 2 solutions by edjones, Theo:
Answer by edjones(8007)

let b=number of beggars and a=amount of money I had before I gave any away.
a-7b=24
a-9b=-32
-------------subtract
2b=56
.
b=28
a=$220.

Answer by Theo(13342)

x = number of beggars
y = money you have.

your formulas are:

y - 7x = 24 (equation 1)
y - 9x = -32 (equation 2)

subtract equation 1 from equation 2 to get:

2x = 56

solve for x to get:

x = 28

substitute in first equation to get:

y - 7*28 = 24

solve for y to get:

y = 7*28 + 24 = 220

you have $220 dollars and there are 28 beggars at your door.

220 - 7*28 = 24 (true)
220 = 9*28 = -32 (true)